# Non-Gaussian enhancement of precision in quantum batteries

### v1

@@ Abstract
We investigate the fundamental limits of converting light into useful work, with a focus on the role of quantum resources in energy harvesting processes. Specifically, we analyze how quantum coherence, non-Gaussianity, and entanglement affect the fluctuations in the energy output of bosonic quantum batteries. Our findings reveal potential pathways to enhance the efficiency and stability of energy extraction from quantum batteries, with implications for the development of quantum thermal machines at the nanoscale. Moreover, this work highlights a tangible thermodynamic quantum advantage, demonstrating how quantum effects can be harnessed to improve the performance of practical energy conversion tasks.
@@
Introduction.–While originally motivated by the industrial necessity of optimizing engine performance, thermodynamics produced some of science’s deepest and most interconnected concepts, ranging from cosmology to information theory to biology Goldstein and Goldstein (1995). Quantum thermodynamics Binder et al. (2018) has the potential to share the same success as its classical counterparts, in a moment where energetic optimization is a fundamental demand and where quantum technologies are receiving notable attention Auffèves (2022). Quantum batteries are one of the most trending topics in the community Campaioli et al. (2024); Konar et al. (2024); Quach et al. (2023); Camposeo et al. (2025); Shi et al. (2025), demonstrating the promise of an energetic quantum advantage in terms of charging power. While being a concept of fundamental interest, power is not a significant figure of merit for practical scenarios, because of the low energetic scales of quantum systems and the resources required to prepare them Kurman et al. (2025). In addition, while having characterized the many ways quantum systems can store work, studies about the practical extraction and usage of such energy are limited Myers et al. (2022).
@@
The edge of quantum mechanics was never about large energies, as proposed by the recent literature on quantum batteries, but about precision, as demonstrated by quantum metrological experiments Liu et al. (2021); Yin et al. (2023); Jia et al. (2024). Precision is also one of the prominent topics of thermodynamics ever since its statistical formulation. In a future fueled by quantum technologies, it is thus hard to imagine that an array of cold atoms will allow the powerful charging of a laptop or anything similar. However, to supply power to the emerging nanotechnologies, realistically, it will be necessary to take into account quantum fluctuations LaHaye et al. (2004). Such devices require an exact amount of energy to produce useful work, and classical fluctuations risk either destroying them or not charging them at all Ussia et al. (2024). Nanobots are typically charged with external power sources to maintain their movement autonomy, using light pulses, magnetic fields, or acoustic waves Mishra (2012); Giri et al. (2021); Zhou et al. (2021). The natural language to describe this dynamic in the quantum setting is the continuous variables formalism Serafini (2023). In the following, we will adopt a quantum optics nomenclature, however, our framework can be easily translated to describe any bosonic quantum field, such as phonons.
In this manuscript, we show new fundamental findings as a means to practical applications for efficient energy processing. The task we propose, precision energy extraction from quantum light, naturally imposes a hierarchy on the type of operations we can perform, enabling the possibility of demonstrating a quantum advantage in terms of a thermodynamic quantity using entangling non-Gaussian unitaries. We optimize quantum batteries under realistic experimental constraints to maximize the signal-to-noise ratio of the work conversion from an energy harvester. Moreover, our scheme is readily implementable in a well-equipped quantum photonics laboratory.
@@
Bosonic batteries.– A quantum battery is defined as a system whose internal Hamiltonian $H_{0}$ has non-degenerate energy levels [^1] In fact, a more relaxed condition of a partially degenerated spectrum, i.e. such that the eigenvalues satisfy the relation $\epsilon_{k}\leq\epsilon_{k+1}$ as long as the bandwidth $\epsilon_{\text{max}}-\epsilon_{\text{min}}$ is non-zero, is sufficient for the definition of a quantum battery Campaioli et al. (2024)., such that energy can be temporarily stored by performing unitary operations $U$ on an initial thermal state $\rho_{0}$ to prepare the system in some excited state $\rho=U\rho_{0}U^{\dagger}$. The energy stored in the battery during this process will afterwards act as the external source for some other device, referred to in this work as energy harvester, capable of transforming the battery’s potential into useful work (e.g. to power some nanotechnological task Jaliel et al. (2019); Culhane et al. (2022); Meng et al. (2025)). In the following, we will assume the battery to be a free propagating electromagnetic field (i.e., a light pulse), however, the same analysis works for any oscillatory system, such as harmonic excitations in crystals.
As for the energy harvester, examples of these kinds of devices include photo-diodes, piezoelectric crystals, or quantum dots Phillips (2021). To maintain the discussion on a general level, we will omit the technicalities of a specific system and assume that this device is anything that generates an electrical current proportional to the number of excitations of the external field $N=\sum_{i}a_{i}^{\dagger}a_{i}$, where we are using the notation of bosonic ladder operators, and we are discretizing the field frequencies to simplify the analysis. The mean electrical current at the output will be given by the product of the mean number of photons $\langle N\rangle$, the repetition rate of the laser source $f$, and a constant that characterizes the efficiency of the process $\mu$. The scheme is synthesized in figure 1. The study of a more realistic model, where the current depends on the continuous envelope of the frequency, will be the subject of future studies.
@@
For $t>0$ the quantum state of light evolves according to the Hamiltonian $H=H_{0}+V(t)$, where the time-dependent potential $V(t)$ is used to change the number of photons to charge the battery. For simplicity, we will assume a constant potential that is instantaneously turned on at the beginning of the protocol and turned off at the end $V(0)=V(\tau)=0$. The difference between the initial and final current produced by the harvester will thus be proportional to
[…]
where the unitary $U=e^{iV\tau/\hbar}$ and $\rho=U\rho_{0}U^{\dagger}$. The fluctuations of this current are given by
[…]
We assume that the potential $V$ can be written as a polynomial of the ladder operators. The degree of this polynomial dramatically changes the type of states that can be produced and will limit the type of fluctuation generated in the electrical current of the harvester. Namely, in the first order, we have classical fields such as coherent states; in the second order, we have quantum squeezed states, which can suppress the fluctuations associated with certain observables Ono et al. (2017); Shchukin et al. (2009). With more than quadratic potentials, we can generate non-Gaussian states that play a prominent role in different areas of quantum information Walschaers (2021). We will show that these states are fundamental to achieving arbitrary precisions in energy harvesting and unlocking the advantage of entangling operations.
In line with this motivation, the figure-of-merit $\Gamma$ that we will seek to maximize will take the form of a signal-to-noise ratio (SNR), between the mean number of photons injected into the battery through the charging process, $\delta N$, and its uncertainty:
$\Gamma=\frac{\delta N}{\Delta N}.$ (4)
Being interested in the more realistic scenario where energy inputs are limited, we fix the maximum allowed $\delta N$ at some value $\epsilon$.
The signal-to-noise ratio $\Gamma$ is related to the second-order intensity correlation function $g^{(2)}(0)$, which is a measure of temporal photon coincidences Glauber (1963)
$g^{(2)}(0)=\bigg{(}\Gamma+\frac{N_{0}}{\Delta N}\bigg{)}^{-2}-\frac{1}{\langle N\rangle}+1.$ (5)
A value of $g^{(2)}(0)$ smaller than unity is a necessary and sufficient condition for the identification of quantum states of light undergoing the phenomenon of antibunching Lemonde et al. (2014). Using Eq.5, this criterion for antibunching can be equivalently rewritten as:
$\bigg{(}\Gamma+\frac{N_{0}}{\Delta N}\bigg{)}^{2}>\langle N\rangle\iff\text{antibunching}$ (6)
Depending on whether the mean photon number of a light source is smaller, equal or greater than its squared variance, their distributions are referred to as super-Poissonian, Poissonian, or sub-Poissonian, respectively. Both Poissonian and super-Poissonian light can be described by a semi-classical theory, in which the light source is modeled as an electromagnetic wave. In contrast, sub-Poissonian light sources are non-classical and are preferable in various applications. They exhibit lower photon-number fluctuations, which is advantageous in high-precision measurements and information processing tasks where minimization of the noise plays a significant role Ann et al. (2019); Berchera and Degiovanni (2019).
@@
Gaussian limitations.– States that can be engineered through Gaussian unitary operations have many desirable properties. They are easy to implement with a high degree of experimental control in state-of-the-art quantum photonics hardware Brask (2022); Zhang et al. (2022). From a theoretical perspective, the properties and dynamics of Gaussian states can be efficiently simulated with a classical computer. Although these states are infinite-dimensional, they allow for a simple mathematical description through symplectic geometry. Despite their relevance from both a fundamental and practical point of view, Gaussian unitaries are limited. In particular, in the context of quantum batteries, we show that they are suboptimal resources for the charging process. For the sake of simplicity, we will initially restrict to the single-mode scenario, in which the initial constraint $\epsilon$ on the battery’s maximum mean photon number reads $\delta N\leq\epsilon$. The generalization to multiple modes is discussed at the end of the Supplementary Material (SM).
The harvested energy signal-to-noise ratio of a single-mode Gaussian state $\rho$ reads:
@@
$\Gamma(\rho)=\frac{N(\rho)-N(\rho_{0})}{\Delta N(\rho)}=\frac{\frac{1}{4}(\operatorname{Tr}(\sigma)-2)+\lVert\bm{\alpha}\rVert^{2}-N(\rho_{0})}{\sqrt{\bm{\alpha}^{T}\sigma\bm{\alpha}+\frac{1}{8}[\operatorname{Tr}(\sigma^{2})-2]}}$ (7)
@@
where $\bm{\alpha}=(\alpha_{1},\alpha_{2})^{T}=\frac{1}{\sqrt{2}}(\langle x\rangle,\langle p\rangle)^{T}$ is the vector of displacements, $\sigma$ is the covariance matrix of state $\rho$, and $\rho_{0}$ is the passive thermal state associated to $\rho$, whose energy is a constant term $N_{0}$ only dependent on its thermal fluctuations parameter $\nu$:
$N_{0}=\frac{1}{2}(\nu-1)\quad\quad\text{with}\quad\quad\nu=\coth\bigg(\frac{\omega}{k_{B}T}\bigg{missing})$ (8)
The explicit expressions for the numerator and denominator of the quotient in Eq.7 read:
Note that given a certain fluctuation degree of our system $\nu$, the ratio $\Gamma$ of any Gaussian state is determined by the squeezing and displacement transformations applied (parametrized by $z$ and $\bm{\alpha}$, respectively), showing no phase dependence.
Let us first consider coherent states, a specific case of Gaussian states that are present in many experimental realizations due to their ease of generation and manipulation. These are the states that most closely resemble the classical wave-like behavior, exhibiting super-Poissonian or Poissonian distributions with $g^{(2)}(0)\geq 1$ (with equality only in the pure case), which cannot display the phenomenon of antibunching. This limitation of coherent states can also be manifest by computing our figure of merit $\Gamma$ upon the substitution $z=1$ (no squeezing is applied, just displacement transformations are performed).
$\Gamma(\rho_{\text{coherent}})=\frac{\lVert\bm{\alpha}\rVert^{2}}{\sqrt{\frac{1}{4}(\nu^{2}-1)+\nu\lVert\bm{\alpha}\rVert^{2}}},$ (11)
The condition $\lVert\bm{\alpha}\rVert^{2}=\delta N\leq\epsilon$ implies that the signal-to-noise ratio for coherent states must be bounded:
$\Gamma(\rho_{\text{coherent}})\leq\lVert\bm{\alpha}\rVert\leq\sqrt{\epsilon}$ (12)
As a consequence, introducing a higher degree of non-classicality is essential to obtain better precision performances in the charging protocol.
Remaining within the Gaussian realm, we now consider the application of quadrature squeezing $z$ to an initial thermal state. Squeezing is an inherently quantum resource that can result in Gaussian states exhibiting sub-Poissonian statistics ($g^{(2)}(0)<1$). In the context of quantum batteries Centrone et al. (2023), if a vacuum or thermal state undergoes squeezing transformations only, its signal-to-noise ratio will again be upper-bounded:
$\Gamma(\rho_{\text{squeezed}})\leq 1$ (13)
@@
$\underbrace{\frac{1}{4}\nu\bigg{(}z+\frac{1}{z}-2\bigg{)}}_{N(\rho)-N(\rho_{0})}\leq\underbrace{\sqrt{\frac{1}{8}\nu^{2}\bigg{(}z^{2}+\frac{1}{z^{2}}\bigg{)}-\frac{1}{4}}}_{\Delta N}\quad\quad\forall z\in(0,1]$ (14)
Squeezed-only batteries perform worse than coherent states if the state contains more than one photon. The optimal Gaussian $\Gamma$ at a certain temperature and given some fixed ergotropic constraint $\epsilon$ will thus be obtained by a combination of both displacement and squeezing such that the energy uncertainty in the resulting state is minimized. The solution to this optimization problem is obtained through the method of Lagrange multipliers, and further discussed in the SM. In Figure 2 the optimal Gaussian ratio is plotted for a range of temperatures and ergotropy constraint values.
Figure 2: Maximum photon number SNR attainable for Gaussian states as a function of temperature and the maximum energy input allowed in the device.
We observe that by combining squeezing and displacement, one can overcome the classical bounds in Eqs. 12, 13, encountered when these operations are performed individually. In fact, for highly squeezed batteries (which appear to be optimal in most cases), we find that the maximum Gaussian ratio can be approximated by the following analytical expression:
The dashed line in the figure represents the boundary $g^{(2)}(0)=1$ between super- and sub-Poissonian photon number distribution of the light source, that is, between classical and quantum behaviour. Note that, since coherent states have super-Poissonian distributions, the acquisition of sub-Poissonian statistics in the optimal states is due to the application of quadrature squeezing. This shows that the energetic signal-to-noise ratio $\Gamma$ is a witness of non-classicality. We now turn to the exploration of a broader set of CV states, which can provide a significant improvement in performance.
Non-Gaussian states.– At zero temperature, the simplest non-Gaussian state, namely the Fock state $|n\rangle$, has a perfectly well-defined expectation value $n$ of the photon number observable with zero variance, therefore yielding $\Gamma\rightarrow\infty$ and becoming the optimal battery. However, as is highlighted in Fig.3, the situation changes dramatically in realistic scenarios. Subject to thermal fluctuations, the uncertainty in the mean photon number of thermal Fock states increases more rapidly than that of squeezed states, which are more robust to temperature increases. Consequently, a thermal regime exists where Fock states no longer provide a significant advantage in the signal-to-noise ratio compared to the Gaussian case, and it thus becomes necessary to explore alternative protocols for precise energy extraction.
@@
In Friis and Huber (2018), an abstract mathematical construction of bosonic states with minimal variance is described and compared to the optimal Gaussian performance. Here we propose an explicit, physically-inspired ansatz for the construction of non-Gaussian states, namely those with the form of a Gaussian plus successive creation operations:
$\rho_{NG}=(a_{1}^{\dagger})^{m}\underbrace{U_{G}\rho_{0}U_{G}^{\dagger}}_{\rho_{G}}(a_{1})^{m}$ (16)
States engineered following this prescription can be realized experimentally in a photonic setup Roeland et al. (2022) through consecutive application of ladder operators (addition or subtraction). Moreover, all bosonic unitaries can be developed in a power series of ladder operators, and in certain cases, it also allows for resummation of an infinite series (see the cat state example in SM). Hence, this ansatz not only allows us to describe practical non-Gaussian states, but, with small adaptations, it potentially encompasses all states of an infinite-dimensional Hilbert space. In addition, the statistical moments of this kind of state allow for exact calculation by applying Wick’s theorem Walschaers et al. (2017). By using a variational algorithm Stornati et al. (2024), we find optimal values of the Gaussian unitary $U_{G}$ to maximize the SNR.

### v2

# Precision bounds for bosonic quantum batteries

@@ Abstract
We study precision charging in bosonic quantum batteries under a finite-energy constraint, using the signal-to-noise ratio (SNR) of delivered excitations as an operational metric directly tied to the energy measured at a load. At the state level, we derive a classical bound whose violation is equivalent to antibunching and certifies non-classicality, and a Gaussian bound whose violation certifies non-Gaussianity under fixed temperature and energy-input constraints. We identify experimentally accessible non-Gaussian families that surpass this Gaussian bound at finite temperature, thereby establishing non-Gaussianity as a resource for enhanced charging precision. Finally, we introduce a linear photodetection model which, under standard linear-response assumptions, propagates these bounds to the photocurrent level and enables both witnesses to be evaluated solely from electrical statistics. Together, these results provide a realistic route to demonstrating an operational quantum advantage—defined as surpassing classical and Gaussian precision bounds—in a thermodynamically motivated energy-conversion task, with plausible near-term applications to the precision charging of fragile nanoscopic loads.
I I. Introduction
@@
While originally motivated by the industrial need to optimize engine performance, thermodynamics has produced some of science’s deepest and most interconnected concepts, ranging from cosmology to information theory to biology [goldstein1995refrigerator]. Quantum thermodynamics [binder2018thermodynamics] has the potential to share the same success as its classical counterparts, in a moment where energetic optimization is a fundamental demand and where quantum technologies are receiving notable attention [auffeves2022quantum]. Quantum batteries are one of the most trending topics in the community [campaioli2024colloquium, Konar_2024, QUACH20232195, Camposeo_2025, shi2025quantumchargingadvantagemultipartite], demonstrating the promise of an energetic quantum advantage in terms of charging power. While being a concept of fundamental interest, power is not a significant figure of merit for practical scenarios, because of the low energetic scales of quantum systems and the resources required to prepare them [kurman2025quantumcomputationquantumbatteries]. In addition, while quantum systems can store work in many ways, studies of the practical extraction and use of such energy are limited [myers2022quantum].
@@
While much of the recent quantum-battery literature emphasizes large energies or charging powers, the edge of quantum mechanics has historically been associated with precision, as demonstrated by quantum metrological experiments [Liu2021, yin2023experimental, jia2024squeezing]. Precision is also a prominent topic in thermodynamics since its statistical formulation. In a future fueled by quantum technologies, it is hard to imagine an array of cold atoms powering a laptop or anything similar. However, to power emerging nanotechnologies, it will be necessary to account for quantum fluctuations [lahaye2004approaching]. Fragile microscopic devices require sufficiently precise energy delivery to produce useful work, while classical fluctuations risk either damaging the device or failing to charge it [ussia2024unlocking]. Examples of externally powered micro/nano-actuators include systems driven by light pulses, magnetic fields, or acoustic waves [mishra2012review, giri2021brief, zhou2021magnetically]. The natural language to describe these processes in the quantum regime is the continuous-variables (CV) formalism [serafini2023quantum]. In what follows, we adopt the quantum-optics nomenclature, but our framework directly applies to any bosonic field (e.g., phonons).
In this work, we analyze precision charging for bosonic quantum batteries under a finite-energy constraint, using the signal-to-noise ratio (SNR) of delivered excitations as an operational metric tied to what a harvester actually measures. On the theory side, we establish two bounds that promote this metric to witnesses: (i) a classical bound (Theorem A) that is equivalent to antibunching and certifies non-classicality, and (ii) a Gaussian bound (Theorem B) at fixed temperature and energy input, whose violation certifies non-Gaussianity. We then identify explicit, experimentally motivated non-Gaussian families that surpass the Gaussian bound in relevant parameter regimes. This complements fluctuation-minimizing constructions in charging [Friis_2018] and SNR-based analyses in specific models [rinaldi2024reliablequantumadvantagequantum] by giving a finite-energy Gaussian bound directly in the conversion metric and by identifying concrete non-Gaussian resources that can, in principle, be certified from electrical readout statistics under standard linear photodetection assumptions.
From a quantum-information perspective, our bounds endow the resource of non-Gaussianity in continuous-variable systems with a task-oriented witness: any state whose conversion SNR exceeds the Gaussian bound $\Gamma_{G}(\nu,\epsilon)$ provides strictly stronger performance than all Gaussian states under the same constraints. This connects non-Gaussianity—identified as a key resource for continuous-variable quantum information processing—to a thermodynamically motivated energy-conversion task, in close analogy to resource-theoretic approaches to non-Gaussianity and other nonclassical resources in CV systems.
In the remainder of this work, Sec. II introduces the bosonic battery–harvester model and defines the conversion SNR $\Gamma$ under a finite-energy constraint. Sec. III derives the classical and Gaussian precision bounds (Theorems A and B), establishes the finite-energy Gaussian bound $\Gamma_{G}(\nu,\epsilon)$, and analyzes non-Gaussian families and multimode interferometric enhancements. Sec. IV discusses photodiode-based implementations and realistic operating regimes, while Sec. V propagates the bounds through a lossy/noisy detection model and benchmarks the detectable precision against thermodynamic-uncertainty limits. We conclude in Sec. VI with a discussion and outlook.
II II. Bosonic batteries
@@
A quantum battery is defined as a system whose internal Hamiltonian $H_{0}$ has non-degenerate energy levels [^1] In fact, a more relaxed condition of a partially degenerated spectrum, i.e. such that the eigenvalues satisfy the relation $\epsilon_{k}\leq\epsilon_{k+1}$ as long as the bandwidth $\epsilon_{\text{max}}-\epsilon_{\text{min}}$ is non-zero, is sufficient for the definition of a quantum battery [campaioli2024colloquium]., such that energy can be temporarily stored by performing unitary operations $U$ on an initial thermal state $\rho_{0}$ to prepare the system in some excited state $\rho=U\rho_{0}U^{\dagger}$. The energy stored in the battery during this process will afterwards act as the external source for some other device (see Fig.1), referred to in this work as energy harvester, capable of transforming the battery’s potential into useful work (e.g. to power some nanotechnological task [jaliel2019experimental, culhane2022extractable, meng2024quantum]). In the following, we will assume the battery to be a free propagating electromagnetic field (i.e., a light pulse), however, the same analysis works for any oscillatory system, such as harmonic excitations in crystals.
As for the energy harvester, examples include photo-diodes, piezoelectric crystals, or quantum dots [phillips2021energy]. We model the harvester as a device producing an electrical response proportional to the excitation number $N=\sum_{i}a_{i}^{\dagger}a_{i}$ of the external field (with bosonic ladder operators and a discrete mode index for clarity). For pulsed operation at repetition rate $f$, the mean electrical output is proportional to $\langle N\rangle$ (a detailed device model is given later), so optimizing the precision of the delivered energy reduces to controlling the mean and fluctuations of $N$.
Initially ($t<0$), the optical field is in a thermal passive state at temperature $T$ for the free Hamiltonian $H_{0}=\sum_{i}\omega_{i}(n_{i}+\tfrac{1}{2})$, i.e.
@@
For $t>0$ the state evolves under $H(t)=H_{0}+V(t)$, where a charging potential $V(t)$ modifies the photon number to charge the battery. We write the protocol unitary as $U=\mathcal{T}\exp\!\big[-\tfrac{i}{\hbar}\int_{0}^{\tau}V(t)\,dt\big]$ and set $V(0)=V(\tau)=0$ for a finite charging window. In the theoretical analysis we work with the number operator and neglect spectral structure; experimental tests fold the detected spectrum into the effective responsivity (see “Implementability”).
In practice $V(t)$ is implemented by a small set of experimentally standard blocks:
$V_{G}(t)$ $=\hbar\!\left[\xi(t)\,a^{\dagger}+\xi^{*}(t)\,a\right]+\frac{\hbar}{2}\,r(t)\!\left(e^{i\phi}a^{2}+e^{-i\phi}a^{\dagger 2}\right),$ (2)
$V_{\mathrm{NG}}(t)$ $=\hbar K(t)\,(a^{\dagger}a)^{2}\ +\ \hbar\chi^{(2)}(t)\!\left(ab+a^{\dagger}b^{\dagger}\right),$ (3)
where $V_{G}$ generates displacement and squeezing (Gaussian unitaries) and $V_{\mathrm{NG}}$ covers Kerr dynamics and heralded photon addition via an ancilla mode $b$ (conditioning on $b$’s detection implements $(a^{\dagger})^{m}$). In the numerical study, we take the pulses as piecewise-constant, which reproduces the “instantaneous on/off” idealization used in the figures. These blocks suffice to prepare the Gaussian families (coherent/squeezed thermal) and the non-Gaussian families (photon-added, cat/kitten) analyzed below.
The mean increase in excitations and its fluctuations are
@@
with $\rho=U\rho_{0}U^{\dagger}$. Our figure of merit is the signal-to-noise ratio (SNR)
$\Gamma=\frac{\delta N}{\Delta N},$ (5)
evaluated under an energy bound $\delta N\leq\epsilon$. In the next section, we analyze how classicality (and Gaussianity) restrict the attainable values of this ratio, thereby allowing us to identify the presence of genuine quantum resources whenever these limits are surpassed.
III III. Results
III.1 A. Classical bound
We say a state is classical if it admits a non-negative Glauber–Sudarshan $P$ representation; coherent states and their mixtures are classical, whereas sub-Poissonian light (Mandel $Q<0$) is non-classical. In our context, classical mixtures of coherent states have Poissonian or super-Poissonian number statistics, so reducing photon-number fluctuations below the Poisson level (at fixed mean) is a signature of non-classicality. The standard second-order correlation $g^{(2)}(0)$ [PhysRev.130.2529] relates to $\Gamma$ via
$g^{(2)}(0)=\left(\Gamma+\frac{N_{0}}{\Delta N}\right)^{-2}-\frac{1}{\langle N\rangle}+1,$ (6)
and $g^{(2)}(0)<1$ is the necessary and sufficient criterion for antibunching [PhysRevA.90.063824].
Theorem A.
For any classical output state, the following bound holds:
$\left(\Gamma+\frac{N_{0}}{\Delta N}\right)^{\!2}\ \leq\ \langle N\rangle,$ (7)
equivalently $g^{(2)}(0)\geq 1$ via Eq. (6). Hence any observed violation of (7) (equivalently, $g^{(2)}(0)<1$) certifies non-classical (sub-Poissonian) light in our precision metric.
Depending on whether the mean photon number is smaller, equal, or greater than its variance, the statistics are termed sub-Poissonian, Poissonian, or super-Poissonian, respectively. Poissonian and super-Poissonian light admit a semi-classical description, whereas sub-Poissonian sources are non-classical and desirable for precision tasks because they suppress photon-number fluctuations [Ann2019, Berchera_2019].
@@
States that can be engineered through Gaussian unitary operations have many desirable properties. They are easy to implement with a high degree of experimental control in state-of-the-art quantum photonics hardware [brask2022gaussianstatesoperations, Zhang2022]. From a theoretical perspective, the properties and dynamics of Gaussian states can be efficiently simulated with a classical computer, and—despite their infinite dimensionality—admit a compact symplectic description. Nevertheless, Gaussian unitaries are limited. In the context of quantum batteries, we show they are suboptimal resources for precision charging [rinaldi2024reliablequantumadvantagequantum].
For simplicity, we first restrict to the single-mode scenario with an energy-input constraint $\delta N\leq\epsilon$; the multimode generalization is discussed in Section III.C.
The harvested-energy SNR of a single-mode Gaussian state $\rho$ is
@@
$\Gamma(\rho)=\frac{N(\rho)-N(\rho_{0})}{\Delta N(\rho)}=\frac{\frac{1}{4}(\operatorname{Tr}(\sigma)-2)+\lVert\bm{\alpha}\rVert^{2}-N(\rho_{0})}{\sqrt{\bm{\alpha}^{T}\sigma\bm{\alpha}+\frac{1}{8}[\operatorname{Tr}(\sigma^{2})-2]}}\,$, (8)
@@
where $\bm{\alpha}=\frac{1}{\sqrt{2}}(\langle x\rangle,\langle p\rangle)^{T}$ is the displacement, $\sigma$ the covariance matrix, and $\rho_{0}$ the associated passive thermal state with
$N_{0}=\tfrac{1}{2}(\nu-1),\qquad\nu=\coth\!\big(\tfrac{\omega}{k_{B}T}\big).$ (9)
Writing the squeezing by a real parameter $0<z\leq 1$ (so that $z$ and $1/z$ are the quadrature variances in shot-noise units), the numerator and denominator of (8) read
$N(\rho)-N(\rho_{0})=\frac{1}{4}\nu\!\left(z+\frac{1}{z}-2\right)+\lVert\bm{\alpha}\rVert^{2},$ (10)
$\Delta N=\sqrt{\frac{1}{8}\nu^{2}\!\left(z^{2}+\frac{1}{z^{2}}\right)-\frac{1}{4}+\nu\!\left(\alpha_{1}^{2}z+\alpha_{2}^{2}\frac{1}{z}\right)}\,.$ (11)
Thus, for fixed thermal fluctuations $\nu$, $\Gamma$ is fully determined by $(z,\bm{\alpha})$ (no phase dependence beyond aligning $\bm{\alpha}$ to the squeezed/anti-squeezed axes).
For coherent states ($z=1$),
$\Gamma(\rho_{\text{coh}})=\frac{\lVert\bm{\alpha}\rVert^{2}}{\sqrt{\frac{1}{4}(\nu^{2}-1)+\nu\lVert\bm{\alpha}\rVert^{2}}},$ (12)
and the constraint $\lVert\bm{\alpha}\rVert^{2}=\delta N\leq\epsilon$ implies
$\Gamma(\rho_{\text{coh}})\leq\sqrt{\epsilon}.$ (13)
If only squeezing is applied to a thermal seed ($\bm{\alpha}=0$), one finds
$\Gamma(\rho_{\text{sq}})\leq 1,$ (14)
@@
$\underbrace{\tfrac{1}{4}\nu\!\left(z+\tfrac{1}{z}-2\right)}_{N(\rho)-N(\rho_{0})}\leq\underbrace{\sqrt{\tfrac{1}{8}\nu^{2}\!\left(z^{2}+\tfrac{1}{z^{2}}\right)-\tfrac{1}{4}}}_{\Delta N}\qquad\forall z\in(0,\infty).$
Theorem B.
Consider the fixed thermal parameter $\nu$ (temperature $T$) and the energy-input constraint $\delta N\leq\epsilon$. Defining the Gaussian bound as
$\Gamma_{G}(\nu,\epsilon)\;:=\;\max_{\text{Gaussian }\rho}\ \Gamma(\rho)\quad\text{subject to }\delta N(\rho)\leq\epsilon,$ (15)
where $\Gamma(\rho)$ is given by Eq. (8), and all Gaussian optimizations are over displacement, single-mode squeezing and passive interferometers; the latter do not change $\Gamma$ (see Appendix G). Then every Gaussian state obeys
$\Gamma(\rho_{G})\ \leq\ \Gamma_{G}(\nu,\epsilon).$
Consequently, any observation of $\Gamma\!>\!\Gamma_{G}(\nu,\epsilon)\$certifies that the state is non-Gaussian (under the same $(\nu,\epsilon)$ constraints).
Squeezed-only batteries perform worse than coherent states once the state contains more than one photon. The optimal Gaussian SNR at a given temperature and fixed ergotropic constraint $\epsilon$ is therefore obtained by combining displacement and squeezing to minimize $\Delta N$ at fixed $\delta N$. We solve this constrained optimization via Lagrange multipliers (see details in Appendix B). Figure 2 shows $\Gamma_{G}(\nu,\epsilon)$ across temperatures and energy budgets; in the highly squeezed regime an accurate analytical approximation is
$\Gamma_{G}(\nu,\epsilon)\ \approx\ \frac{8\epsilon}{\nu^{2}\!\left[\,3\Big(\frac{4\epsilon}{\nu}+2\Big)^{\!\frac{2}{3}}-2+\Big(\frac{4\epsilon}{\nu}+2\Big)^{\!-\frac{2}{3}}\right]}.$ (16)
Figure 2: Maximum photon-number SNR attainable for Gaussian states, $\Gamma_{G}(\nu,\epsilon)$, as a function of temperature (via $\nu$) and maximum energy input $\epsilon$. The dashed line marks $g^{(2)}(0)=1$ (classical/non-classical boundary).
We observe that by combining squeezing and displacement, one can surpass the individual coherent/squeezed bounds (13)–(14), and yield optimal Gaussian states lying in the sub-Poissonian region ($g^{(2)}(0))<1$. In the next section we explore how genuinely non-Gaussian states can exceed $\Gamma_{G}(\nu,\epsilon)$, thereby realizing a precision advantage beyond the entire Gaussian sector.
III.3 C. Non-Gaussian states
At zero temperature, the simplest non-Gaussian state, namely the Fock state $|n\rangle$, has a perfectly well-defined expectation value $n$ of the photon-number observable with zero variance, therefore yielding $\Gamma\to\infty$ and becoming the optimal battery. However, as shown in Fig. 3, the situation changes dramatically in realistic scenarios. Subject to thermal fluctuations, the uncertainty in the mean photon number of thermal Fock states increases more rapidly than that of squeezed states, which are more robust to temperature increases. Consequently, a thermal regime exists where Fock states no longer provide a significant advantage in the signal-to-noise ratio compared to the Gaussian case, and it thus becomes necessary to explore alternative protocols for precise energy extraction.
@@
In Ref. [Friis_2018], Friis and Huber introduced a mathematical construction of bosonic states that minimize fluctuations and compared them to the optimal Gaussian performance within a charging framework. More recently, Rinaldi et al. studied charging performance using a signal-to-noise figure of merit in a Jaynes–Cummings (qubit–cavity) setting, identifying an advantage offered by pure non-classical resources [rinaldi2024reliablequantumadvantagequantum]. Here we instead derive a finite-energy Gaussian bound $\Gamma_{G}(\nu,\epsilon)$ for the conversion SNR $\Gamma$ (Theorem B) under an explicit energy-input constraint $\delta N\leq\epsilon$, and we exhibit physically motivated, experimentally accessible non-Gaussian families that violate this bound at finite temperature. Unlike Ref. [Friis_2018], which constructs minimum-fluctuation states without an explicit detection model, and Ref. [rinaldi2024reliablequantumadvantagequantum], which focuses on a specific Jaynes–Cummings charging setting, our analysis is tailored to bosonic batteries delivering energy to a bosonic load and links directly to a measurable quantity (the delivered excitation number), with a simple linear photodetection model that makes the Gaussian bound directly testable from photocurrent statistics. In particular, within this framework, any observed $\Gamma>\Gamma_{G}(\nu,\epsilon)$ certifies that the prepared state is non-Gaussian (see Theorem B).
Beyond the restriction to the Gaussian realm, we adopt an explicit, physically inspired ansatz:
$\rho_{\mathrm{NG}}\;=\;(a^{\dagger})^{m}\,\underbrace{U_{G}\rho_{0}U_{G}^{\dagger}}_{\rho_{G}}\,(a)^{m},$ (17)
i.e., photon addition (or subtraction) atop a Gaussian block $U_{G}$ (displacement and squeezing). Exact closed forms for moments via Wick’s theorem are given in Appendix C. States engineered by (17) arise naturally in photonic platforms via conditional operations on weakly tapped beams and single-photon detection, and multi-photon addition/subtraction is now routine [roeland2022mode]. More generally, bosonic unitaries can be expanded in ladder operators, and in some cases resummed (e.g., cat/kitten states; see Appendix D), so this family captures both general and practically-relevant non-Gaussian resources. Importantly, the statistical moments of (17) can be computed exactly by Wick’s theorem [PhysRevA.96.053835], which enables efficient evaluation of $\Gamma$ and gradient-based optimization. All non-Gaussian expectations are computed exactly without Fock-space truncation; in Appendix F we show that naive truncation can spuriously violate the Gaussian bound. Using a variational algorithm [fedeypaolo], we optimize the Gaussian block $U_{G}$ for fixed $m$ and constraints $(\nu,\epsilon)$.